OpenStudy (anonymous):

how do i find the indefinate integral of -sin(x)/cos^2(x)dx?

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits \frac{1}{\cos^2x}\left(-\sin x\;dx\right)\] If I split it up like that, can you see your u-substitution easier?

OpenStudy (anonymous):

i am not sure

OpenStudy (anonymous):

what do i do from there?

zepdrix (zepdrix):

\[\Large\bf\sf u=\cos x\]

zepdrix (zepdrix):

What do you get for `du` ?

OpenStudy (anonymous):

-cos(x)?

zepdrix (zepdrix):

No silly! We take our substitution and then take its derivative with respect to x.\[\Large\bf\sf u=\cos x, \qquad\implies\qquad \frac{du}{dx}=-\sin x\]Then simply "multiply" your dx to the other side to solve for du.\[\Large\bf\sf du=-\sin x\;dx\]

zepdrix (zepdrix):

\[\Large\bf\sf du=(-\sin x\;dx)\]The du is showing up somewhere in our integral, yes?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits\limits \frac{1}{(\color{royalblue}{\cos x})^2}\left(\color{orangered}{-\sin x\;dx}\right)\] \[\Large\bf\sf \color{royalblue}{u=\cos x}, \qquad\qquad\qquad \color{orangered}{du=-\sin x\;dx}\] Do you understand how we'll be plugging in the pieces for u?

OpenStudy (anonymous):

yes i s

OpenStudy (anonymous):

yes i see it

zepdrix (zepdrix):

Ok good. \[\Large\bf\sf \int\limits\limits\limits \frac{1}{(\color{royalblue}{\cos x})^2}\left(\color{orangered}{-\sin x\;dx}\right)\quad\to\quad \int\limits\limits\limits \frac{1}{(\color{royalblue}{u})^2}\left(\color{orangered}{du}\right)\]

zepdrix (zepdrix):

From here we can apply a rule of exponents to write it like this,\[\Large\bf\sf \int\limits u^{-2}\;du\]And then just apply your power rule for integration.

OpenStudy (shamil98):

Very nice job at explaining this @zepdrix! Keep it up :)

zepdrix (zepdrix):

color ftw ^^

OpenStudy (anonymous):

okay thank you, i think i got it from here, :)

zepdrix (zepdrix):

Cool c: